# Influence of Modeling Assumptions on the Inferred Dynamical State of Resonant Systems: A Case Study of the HD 45364 System

### v1

@@ Abstract
Planetary systems exhibiting mean-motion resonances (MMRs) offer unique opportunities to study the imprint of disk-induced migration on the orbital architectures of planetary systems. The HD 45364 system, discovered via the radial velocity (RV) method to host two giant planets in a 3:2 MMR, has been the subject of several studies attempting to reconstruct the system’s orbital migration history based on its present-day resonant configuration. Recently, Li et al. (2022) called into question the system’s residence in the 3:2 MMR based on a revised orbital solution derived from an expanded set of RV observations that extend the time baseline of the original discovery data by over a decade. However, we show that inferences about the planets’ dynamical state with respect to the 3:2 MMR are sensitive to the particular prior assumptions adopted in the orbital modeling. Using $N$-body dynamical models, we show that orbital solutions constrained to reside deep in the 3:2 MMR fit the RV data with a similar quality to unconstrained orbital solutions. We conclude that the RV observations of HD 45364 are consistent with orbital configurations produced by smooth migration and resonance capture. We further show that past convergent orbital migration can reproduce the system’s present-day orbital configuration provided that the ratio of migration to eccentricity damping timescales, $K$, was in the range $10\lesssim K\lesssim 175$. We also find that dynamical interactions in the system can break the usual mass-inclination degeneracy inherent to Keplerian models of RV observations and constrain the planets’ absolute masses to within a factor of $\sim 2$.
Exoplanets (498) – Orbital resonances (1181) – Celestial mechanics (211) – Radial velocity (1332)
journal: ApJsoftware: REBOUND (Rein & Liu, 2012), emcee (Foreman-Mackey et al., 2013), REBOUNDx (Tamayo et al., 2020)
@@ 1 Introduction
Numerous pairs of giant planets in mean-motion resonances (MMRs) have been discovered via the radial velocity (RV) method (e.g. Mayor et al., 2004; Lee et al., 2006; Tinney et al., 2006; Correia et al., 2009; Niedzielski et al., 2009; Johnson et al., 2010, 2011; Wright et al., 2011; Robertson et al., 2012; Wittenmyer et al., 2013, 2016; Giguere et al., 2015; Luque et al., 2019; Trifonov et al., 2019). The resonant orbital configurations of these systems are generally interpreted as products of past convergent orbital migration induced by planet interactions with the protoplanetary disk (Goldreich & Tremaine, 1980), as such migration generically causes planets to capture into any low-order MMRs they encounter. Thus, systems hosting resonant planets present attractive choices for studying orbital migration and its role in planetary formation.
One such system is HD 45364, which consists of a K0V star with mass of $0.82~{}{M}_{\odot}$ and two approximately Jupiter-mass planets with orbits near a 3:2 commensurability, discovered by Correia et al. (2009) using the High Accuracy Radial Velocity Planet Searcher (HARPS; Mayor et al., 2003). The original best-fit radial velocity solution of Correia et al. (2009) indicated that the two planets were on slightly eccentric orbits and experienced large librations within the $3$:$2$ MMR, with $\sim 70\degree$ excursions of the resonant angles away from their equilibrium values. Subsequent work by Rein et al. (2010) used hydrodynamical simulations of planet-disk interactions to find potential migration histories that could produce orbital configurations that were consistent with the observed RV data. They found orbital solutions arising from their migration simulations that differed significantly from the best-fit solution reported by Correia et al. (2009) but which reproduced the observed data with approximately equal statistical significance. Similar conclusions were reached by Correa-Otto et al. (2013) using $N$-body simulations with parameterized migration and eccentricity damping forces. Finally, Hadden & Payne (2020) showed, using a Bayesian model comparison framework, that the available RV data for HD 45364 were fully consistent with the planet pair residing in a zero-libration amplitude resonant configuration that would naturally arise under the effects of smooth migration and eccentricity damping forces.
@@ 2 Radial Velocity Fitting
The original RV observations of HD 45364 used by Correia et al. (2009) were taken by the HARPS telescope over a period of approximately $1600$ days from December $2003$ to April $2008$. Further observations of the system by HARPS have since extended its observational baseline to September $2017$, spanning over $5000$ days in total. Trifonov et al. (2020) improved the HARPS RV precision by recomputing the RVs from observed spectra using the Spectrum Radial Velocity Analyser pipeline of Zechmeister et al. (2018) and correcting a number of systematic errors. The new HARPS data were later published as part of the HARPS-RVBank archive (Trifonov et al., 2020), which we use here. In addition, an upgrade to HARPS’s optical fibres in May $2015$ (Lo Curto et al., 2015) changed the instrumental profile and thus the RV offset between the pre- and post-upgrade RV observations (Trifonov et al., 2020). As such, we treat the pre- and post-upgrade RV data as taken from two separate instruments in this work. Finally, we also use $7$ RV observations taken by HIRES from December $2009$ to September $2021$, provided in Table $1$ of Li et al. (2022).
Here we report our fits to the full set of RV data available for HD 45634, spanning a total observational baseline of $\sim 6500$ days. We begin in Section 2.1 by describing our $N$-body dynamical model, which we fit under relatively uninformative priors on the planets’ orbital parameters. In Section 2.2, we explore how the inferred dynamical state of the system with respect to the 3:2 MMR is influenced by the model priors and assumptions about the level of measurement noise.
@@ 2.1 $N$-body dynamical model
We fit the RV observations of HD 45364 using an $N$-body model that accounts for Newtonian gravitational interactions among the star and two planets in order to compute the star’s radial velocity. Integrations for this dynamical model and all other simulations presented in this paper are conducted with the REBOUND (Rein & Liu, 2012) code for numerical $N$-body integration, using the IAS15 integrator (Rein & Spiegel, 2015). We fit our model to three sets of RV observations: the pre- and post-upgrade HARPS observations (which we denote HARPS1 and HARPS2 respectively), and the HIRES observations.
The planets’ orbits are parameterized in terms of their orbital periods, $P_{i}$, eccentricities, $e_{i}$, times of conjunction, $T_{i}$, and arguments of pericenter, $\omega_{i}$, all specified at a reference epoch of $t_{0}=54422.79+2.4\times 10^{6}$ BJD. We parameterize the planets’ masses, $m_{i}$, in our model as functions of the planets’ orbital parameters and RV semi-amplitudes, $K_{i}$, according to
@@
where $M_{*}$ is the host star mass (which we fix to be $M_{*}=0.82\,M_{\odot}$), and $G$ is Newton’s universal gravitational constant. We assume the planets’ orbits to be coplanar but allow their orbital planes to be inclined with respect to the sky plane by an angle, $I$, and include $\sin(I)$ as a free parameter in our $N$-body model. Our model also includes radial velocity offsets, $\gamma_{j}$, for the three sets of observations HARPS1, HARPS2, and HIRES. Finally, our model includes three instrumental jitter parameters $\sigma_{\mathrm{jit},j}$, one for each set of observations. Writing $\boldsymbol{\theta}_{\mathrm{orb}}=\{\sin I\}\cup\{K_{i},P_{i},e_{i},T_{i},\omega_{i}\}_{i=1}^{2}$ to denote the subset model parameters governing the orbital data, the log-likelihood of a set of model parameters can be written as
@@
$-\sum_{i=1}^{N}\,\ln\sqrt{2\pi\left(\sigma_{j,i}^{2}+\sigma_{\mathrm{jit},j}\right)}$ (2)
where the index $j$ runs over the three sets of radial velocity data from HARPS1, HARPS2, and HIRES, $v_{j,i}$ denotes $i$th observed radial velocity from the $j$th set of observations, $\sigma_{j,i}$ denotes this velocity’s reported measurement uncertainty, and $\bar{v}_{j,i}(\boldsymbol{\theta}_{\mathrm{orb}})$ denotes the star’s radial velocity predicted by our $N$-body dynamical model at the time of observation.
We use Markov Chain Monte Carlo (MCMC) sampling to estimate the posterior distribution of our model parameters. We adopt priors that are uniform on the interval $[0,\infty)$ for planets’ orbital periods, $P_{i}$, semi-amplitudes, $K_{i}$, and times of conjunction, $T_{i}$. Our prior on the sine of the planets’ orbital inclination, $\sin I$, is uniform between 0 and 1. We adopt the standard procedure of parameterizing planets’ orbital eccentricities, $e_{i}$, and arguments of pericenter, $\omega_{i}$, by the variables $\sqrt{e_{i}}\cos\omega_{i}$ and $\sqrt{e_{i}}\sin\omega_{i}$ when sampling, while imposing a prior that planets’ eccentricities be less than 1. Finally, our priors on the value of the instrumental jitters, $\sigma_{\mathrm{jit},j}$, are uniform between 0 and 10 m/s. The MCMC sampling is carried out using the emcee (Goodman & Weare, 2010; Foreman-Mackey et al., 2013) software package. We initialize $50$ chains in a multivariate Gaussian distribution centered on the best-fit parameters obtained by minimizing the log-likelihood. The chains are then evolved for $50,000$ steps, generating a total of $2.5$ million posterior samples. The results for our RV fit and MCMC sampling are summarized in Figure 1. The top panel shows the $1\sigma$, $2\sigma$ and $3\sigma$ uncertainties of our fit RV curve together with the observed data. We obtain best-fit instrumental jitter values of $2.15$ m/s, $1.49$ m/s and $4.43$ m/s for the HARPS1111, HARPS2222 and HIRES telescopes, respectively.
@@
$m_{b}\,\left[M_{J}\right]$ $0.198$ $0.199$ $0.256$ $0.470$
$P_{b}$ [days] $227.86$ $226.89$ $228.01$ $229.23$
$e_{b}$ $0.0457$ $0.0181$ $0.0638$ $0.120$
$T_{b}$ [BJD - $2.4\times 10^{6}$ days] $53938.67$ $53933.73$ $53941.19$ $53949.03$
$\omega_{b}$ [rad] $-3.11$ $-2.99$ $-2.27$ $2.94$
$m_{c}\,\left[M_{J}\right]$ $0.576$ $0.588$ $0.746$ $1.38$
$P_{c}$ [days] $344.03$ $343.15$ $343.91$ $344.67$
$e_{c}$ $0.0146$ $0.00212$ $0.0220$ $0.0548$
$T_{c}$ [BJD - $2.4\times 10^{6}$ days] $54705.92$ $54701.06$ $54705.58$ $54709.17$
$\omega_{c}$ [rad] $1.25$ $-2.00$ $1.00$ $2.51$
$\sin\left(I\right)$ $1.00$ $0.42$ $0.77$ $0.98$
$\sigma_{\mathrm{jit},1}$ [m/s] $2.15$ $1.85$ $2.43$ $2.99$
$\sigma_{\mathrm{jit},2}$ [m/s] $1.49$ $1.01$ $2.31$ $4.29$
$\sigma_{\mathrm{jit},3}$ [m/s] $4.43$ $2.69$ $6.39$ $9.54$
Figure 1: Summary of radial velocity fitting results and uncertainties from the MCMC posterior distributions for HD 45364, using data from the pre- and post-upgrade HARPS (HARPS1 and HARPS2, respectively) as well as the HIRES instruments. Top: Our fit RV signal (red) is plotted over the observed RVs, with $1\sigma$, $2\sigma$ and $3\sigma$ uncertainties illustrated by the shaded regions. The HARPS data points to the right of the black line are new data obtained from the HARPS-RVBank archive (Trifonov et al., 2020) since the original analysis of Correia et al. (2009). Time is given as Barycentric Julian Date (BJD) - $2.4\times 10^{6}$ days. Bottom: Table showing the log-likelihood and parameter values for the best-fit solution obtained by maximizing the log-likelihood, as well as the $5\%$, $50\%$ (median) and $95\%$ quantiles of the $1$-D marginalized posterior distributions obtained from MCMC sampling.
@@
Figure 2: The marginal $2$-D MCMC posterior distributions of $\sin\left(I\right)$ against planet masses $m_{b}$ (left) and $m_{c}$ (right), with the planets assumed to be in coplanar orbits. The contours show the $1\sigma$, $2\sigma$ and $3\sigma$ bounds of the posterior distribution. The points corresponding to the best-fit solution obtained by maximizing the log-likelihood $\ln\mathcal{L}$ as given in Table 1 are marked in red.
@@ 2.2 The influence of model priors and measurement noise on resonant dynamics
The solid lines in Figure 3 show the cumulative distribution for the maximum deviation of the planets’ resonant angles from equilibrium, measured from short $N$-body integrations spanning $500$ orbits of the inner planet for a random selection of $50,000$ samples from our full MCMC posterior. For each posterior sample, the resonant angles $\theta_{b}=3\lambda_{c}-2\lambda_{b}-\varpi_{b}$ and $\theta_{c}=3\lambda_{c}-2\lambda_{b}-\varpi_{c}$ were recorded at $5000$ uniformly spaced times in the integration and the maximum deviation from their respective equilibrium values of 00 and $\pi$ were computed. Figure 3 clearly indicates that most of the posterior samples have large libration amplitudes, with roughly half of the posterior samples showing libration amplitudes greater than $90^{\circ}$ in both resonant angles and a majority of posterior samples ($\gtrsim 60\%$) showing circulation in angle $\theta_{c}$. These results nominally support the conclusion of Li et al. (2022) that the HD 45364 planets are not locked in the 3:2 MMR.
Figure 3: The cumulative distribution of the maximum deviation from equilibrium for the resonant angles $\theta_{b}$ (blue) and $\theta_{c}$ (orange), measured over short $N$-body integrations, of $50,000$ samples drawn randomly from the MCMC posteriors for the $N$-body dynamical model fit under uninformative priors described in Section 2.1 (solid lines) and the model described in Section 2.2 fit under a resonance-weighted prior (i.e. weighted towards small libration amplitudes), with $S=0.1$ (dashed lines). A large fraction of orbital configurations in the uninformative prior model’s MCMC posteriors circulate or have large libration amplitudes ($>90\degree$). In contrast, virtually all resonant angles in the resonance-weighted prior model’s MCMC posteriors have small libration amplitudes ($<90\degree$), showing that the chosen model priors strongly affect the inferred libration amplitudes of the system.
However, Jensen & Millholland (2022) demonstrate that measurement noise can cause a systematic bias to resonant libration amplitudes inferred from radial velocities. This bias is essentially caused by the fact that higher libration amplitude dynamical configurations occupy a much larger phase space volume. Consequently, under model priors that are relatively uniform over phase space, noisier data will lead to broader posterior distributions that, in turn, become increasingly skewed towards large libration amplitudes or circulating resonant angles. In this section, we explore the sensitivity of the inferred resonant libration amplitudes of the HD 45364 planets to assumptions about measurement noise and model priors.
@@
where $\mathcal{L}$ is the likelihood function of our uninformative prior RV model given in Equation (2) and $\hat{\sigma}_{\mathrm{jit},j}$ are the best-fit jitter values determined via a maximum-likelihood fit to the RV data using the uninformative prior RV model. In Figure 4, we numerically maximize the modified log-likelihood, $\ln\mathcal{L}^{\prime}$, on a grid of fixed values for the parameters $f$ and $S$. We then plot contour levels of $\ln\mathcal{L}$, i.e., the un-modified, uninformative RV model log-likelihood, evaluated at the numerically determined parameters that maximize the modified likelihood, $\mathcal{L}^{\prime}$. The $n\sigma$ confidence regions are estimated as $\Delta\ln\mathcal{L}=-\frac{1}{2}n^{2}$, where $\Delta\ln\mathcal{L}$ is the difference between the log-likelihood function and its global maximum value. If the data strongly favored dynamical configurations in which the resonant angles circulated or possessed large libration amplitudes, we would expect to find that the values of $\ln\mathcal{L}$ decrease significantly for sufficiently small values of $S$. However, this is not what we see in Figure 4. Instead, contours of constant likelihood are nearly independent of the assumed value of $S$ except at the lowest values of the jitter scaling factor, $f$. Thus, we conclude that the prevalence of dynamical configurations with circulating and large libration amplitude resonant angles found among the posterior samples of our uninformative prior RV model is the result of these dynamical configurations occupying larger prior volume, rather than the data disfavoring low libration amplitude configurations. Furthermore, the fact that low libration amplitude solutions become disfavored relative to higher amplitude solutions for small values of $f$ indicates that underestimated measurement noise can lead one to improperly infer large libration amplitudes.

### v2

@@ Abstract
Planetary systems exhibiting mean-motion resonances (MMRs) offer unique opportunities to study the imprint of disk-induced migration on the orbital architectures of planetary systems. The HD 45364 system, discovered via the radial velocity (RV) method to host two giant planets in a 3:2 MMR, has been the subject of several studies attempting to reconstruct the system’s orbital migration history based on its present-day resonant configuration. Recently, Li et al. (2022) called into question the system’s residence in the 3:2 MMR based on a revised orbital solution derived from an expanded set of RV observations that extend the time baseline of the original discovery data by over a decade. However, we show that inferences about the planets’ dynamical state with respect to the 3:2 MMR are sensitive to the particular prior assumptions adopted in the orbital modeling. Using $N$-body dynamical models, we show that orbital solutions constrained to reside deep in the 3:2 MMR fit the RV data with a similar quality to unconstrained orbital solutions. We conclude that the RV observations of HD 45364 are consistent with orbital configurations produced by smooth migration and resonance capture. We further show that past convergent orbital migration can reproduce the system’s present-day orbital configuration provided that the ratio of migration to eccentricity damping timescales, $K$, was in the range $11\lesssim K\lesssim 144$. We also find that dynamical interactions in the system can break the usual mass-inclination degeneracy inherent to Keplerian models of RV observations and constrain the planets’ absolute masses to within a factor of $\sim 1.5$.
Exoplanets (498) – Orbital resonances (1181) – Exoplanet dynamics (490) – Celestial mechanics (211) – Radial velocity (1332)
software: REBOUND (Rein & Liu, 2012), emcee (Foreman-Mackey et al., 2013), REBOUNDx (Tamayo et al., 2020)
@@ 1 Introduction
Numerous pairs of giant planets in mean-motion resonances (MMRs) have been discovered via the radial velocity (RV) method (e.g. Mayor et al., 2004; Lee et al., 2006; Tinney et al., 2006; Correia et al., 2009; Niedzielski et al., 2009; Johnson et al., 2010, 2011; Wright et al., 2011; Robertson et al., 2012; Wittenmyer et al., 2013, 2016; Giguere et al., 2015; Luque et al., 2019; Trifonov et al., 2019). The resonant orbital configurations of these systems are generally interpreted as products of past convergent orbital migration induced by planet interactions with the protoplanetary disk (Goldreich & Tremaine, 1980), as such migration generically causes planets to be captured into any low-order MMRs they encounter. Thus, systems hosting resonant planets present attractive choices for studying orbital migration and its role in planetary formation.
One such system is HD 45364, which consists of a K0V star with mass of $0.82~{}{M}_{\odot}$ and two approximately Jupiter-mass planets with orbits near a 3:2 commensurability, discovered by Correia et al. (2009) using the High Accuracy Radial Velocity Planet Searcher (HARPS; Mayor et al., 2003). The original best-fit RV solution of Correia et al. (2009) indicated that the two planets were on slightly eccentric orbits and experienced large librations within the $3$:$2$ MMR, with $\sim 70\degree$ excursions of the resonant angles away from their equilibrium values. Subsequent work by Rein et al. (2010) used hydrodynamical simulations of planet-disk interactions to find potential migration histories that could produce orbital configurations that were consistent with the observed RV data. They found orbital solutions arising from their migration simulations that differed significantly from the best-fit solution reported by Correia et al. (2009) but which reproduced the observed data with approximately equal statistical significance. Similar conclusions were reached by Correa-Otto et al. (2013) using $N$-body simulations with parameterized migration and eccentricity damping forces. Finally, Hadden & Payne (2020) showed, using a Bayesian model comparison framework, that the available RV data for HD 45364 were fully consistent with the planet pair residing in a zero-libration amplitude resonant configuration that would naturally arise under the effects of smooth migration and eccentricity damping forces.
@@ 2 Radial Velocity Fitting
The original RV observations of HD 45364 used by Correia et al. (2009) were taken by the HARPS telescope over a period of approximately $1600$ days from December $2003$ to April $2008$. Further observations of the system by HARPS have since extended its observational baseline to September $2017$, spanning over $5000$ days in total. Trifonov et al. (2020) improved the HARPS RV precision by recomputing the RVs from observed spectra using the Spectrum Radial Velocity Analyser pipeline of Zechmeister et al. (2018) and correcting a number of systematic errors. The new HARPS data were later published as part of the HARPS-RVBank archive111github.com/3fon3fonov/HARPS_RVBank(Trifonov et al., 2020), totalling $122$ RV observations. In addition, an upgrade to HARPS’s optical fibres in May $2015$ (Lo Curto et al., 2015) changed the instrumental profile and thus the RV offset between the pre- and post-upgrade RV observations (Trifonov et al., 2020). As such, we treat the pre- and post-upgrade RV data as taken from two separate instruments in this work. Finally, we also use $7$ RV observations taken by HIRES from December $2009$ to September $2021$, provided in Table $1$ of Li et al. (2022).
Here we report our fits to the full set of RV data available for HD 45634, consisting of $129$ data points spanning a total observational baseline of $\sim 6500$ days. [^2] We note that Li et al. (2022) use only $121$ total data points ($114$ from HARPS and $7$ from HIRES) in their analysis, as in two separate cases, they combine five HARPS observations taken over a short time span into a single observation. We begin in Section 2.1 by describing our $N$-body dynamical model, which we fit under relatively uninformative priors on the planets’ orbital parameters. In Section 2.2, we explore how the inferred dynamical state of the system with respect to the 3:2 MMR is influenced by the model priors and assumptions about the level of measurement noise.
@@ 2.1 $N$-body dynamical model
We fit the RV observations of HD 45364 using an $N$-body model that accounts for Newtonian gravitational interactions among the star and two planets in order to compute the star’s RV. Integrations for this dynamical model and all other simulations presented in this paper are conducted with the REBOUND (Rein & Liu, 2012) code for numerical $N$-body integration, using the IAS15 integrator (Rein & Spiegel, 2015) for the integration and Jacobi elements for the coordinate frame. We fit our model to three sets of RV observations: the pre- and post-upgrade HARPS observations (which we denote HARPS1 and HARPS2 respectively), and the HIRES observations.
The planets’ orbits are parameterized in terms of their orbital periods, $P_{i}$, eccentricities, $e_{i}$, times of conjunction, $T_{i}$, and arguments of pericenter, $\omega_{i}$, all specified at a reference epoch of $t_{0}=54422.79+2.4\times 10^{6}$ BJD. This value is chosen as the median time of the RV observations for the system. We parameterize the planets’ masses, $m_{i}$, in our model as functions of the planets’ orbital parameters and RV semi-amplitudes, $K_{i}$, according to
@@
where $M_{*}$ is the host star mass (which we fix to be $M_{*}=0.82\,M_{\odot}$), and $G$ is Newton’s universal gravitational constant. We assume the planets’ orbits to be coplanar but allow their orbital planes to be inclined with respect to the sky plane by an angle, $I$, and include $\sin(I)$ as a free parameter in our $N$-body model. Our model also includes RV offsets, $\gamma_{j}$, for the three sets of observations HARPS1, HARPS2, and HIRES. Finally, our model includes three instrumental jitter parameters $\sigma_{\mathrm{jit},j}$, one for each set of observations. Writing $\boldsymbol{\theta}_{\mathrm{orb}}=\{\sin I\}\cup\{K_{i},P_{i},e_{i},T_{i},\omega_{i}\}_{i=1}^{2}$ to denote the subset model parameters governing the orbital data, the log-likelihood of a set of model parameters can be written as
@@
$-\sum_{i=1}^{N_{j}}\,\ln\sqrt{2\pi\left(\sigma_{j,i}^{2}+\sigma_{\mathrm{jit},j}^{2}\right)}$ (2)
where the index $j$ runs over the three sets of RV data from HARPS1, HARPS2, and HIRES, $v_{j,i}$ denotes $i$th observed RV from the $j$th set of observations, $\sigma_{j,i}$ denotes this velocity’s reported measurement uncertainty, and $\bar{v}_{j,i}(\boldsymbol{\theta}_{\mathrm{orb}})$ denotes the star’s RV predicted by our $N$-body dynamical model at the time of observation.
We use Markov Chain Monte Carlo (MCMC) sampling to estimate the posterior distribution of our model parameters. We adopt priors that are uniform on the interval $[0,\infty)$ for planets’ orbital periods, $P_{i}$, semi-amplitudes, $K_{i}$, and times of conjunction, $T_{i}$. Our prior on the sine of the planets’ orbital inclination, $\sin I$, is uniform between 0 and 1. We adopt the standard procedure of parameterizing planets’ orbital eccentricities, $e_{i}$, and arguments of pericenter, $\omega_{i}$, by the variables $\sqrt{e_{i}}\cos\omega_{i}$ and $\sqrt{e_{i}}\sin\omega_{i}$ when sampling, while imposing a prior that planets’ eccentricities be less than 1. Finally, our priors on the value of the instrumental jitters, $\sigma_{\mathrm{jit},j}$, are uniform between 0 and 10 m s-1. The MCMC sampling is carried out using the emcee (Goodman & Weare, 2010; Foreman-Mackey et al., 2013) software package. We initialize $50$ chains in a multivariate Gaussian distribution centered on the best-fit parameters obtained by minimizing the log-likelihood. The chains are then evolved for $50,000$ steps, generating a total of $2.5$ million posterior samples. We compute the autocorrelation lengths of our MCMC chains for each model parameter and estimate that our simulations produce $\gtrsim 3000$ independent posterior samples by dividing the total number of samples by the longest computed autocorrelation length.
The resulting RV fit is shown in Figure 1. The top panel shows the $1\sigma$, $2\sigma$ and $3\sigma$ uncertainties of our fit RV curve together with the observed data. We obtain best-fit instrumental jitter values of $1.41$, $0.694$ and $3.00$ m s-1 for the HARPS1111, HARPS2222 and HIRES telescopes, respectively. The maximum log-likelihood solution and the $5\%$, $50\%$ and $95\%$ quantiles of the $1$D posterior distributions obtained from MCMC sampling are reported in Table 1.
Table 1: Maximum Log-Likelihood Fit and MCMC Posterior Quantiles for Fitted Parameters
@@
$m_{b}\,\left[M_{J}\right]$ $0.191$ $0.191$ $0.222$ $0.320$
$P_{b}$ [days] $227.88$ $227.20$ $227.93$ $228.69$
$e_{b}$ $0.0505$ $0.0328$ $0.0613$ $0.0917$
$T_{b}$ [BJD - $2.4\times 10^{6}$ days] $52799.70$ $52793.08$ $52801.02$ $52808.58$
$\omega_{b}$ [rad] $-2.95$ $-3.03$ $-2.51$ $2.94$
$m_{c}\,\left[M_{J}\right]$ $0.550$ $0.557$ $0.641$ $0.932$
$P_{c}$ [days] $344.06$ $343.53$ $344.01$ $344.49$
$e_{c}$ $0.00991$ $0.00132$ $0.0136$ $0.0353$
$T_{c}$ [BJD - $2.4\times 10^{6}$ days] $52985.87$ $52982.85$ $52986.16$ $52989.33$
$\omega_{c}$ [rad] $1.34$ $-1.99$ $1.03$ $2.51$
$\sin\left(I\right)$ $1.00$ $0.594$ $0.860$ $0.988$
$\sigma_{\mathrm{jit},\mathrm{HARPS1}}$ [m/s] $1.41$ $1.19$ $1.50$ $1.76$
$\sigma_{\mathrm{jit},\mathrm{HARPS2}}$ [m/s] $0.694$ $0.17$ $0.96$ $1.66$
$\sigma_{\mathrm{jit},\mathrm{HIRES}}$ [m/s] $3.00$ $1.89$ $4.00$ $7.57$
Note. The orbital periods $P_{i}$, times of conjunction $T_{i}$, eccentricities $e_{i}$ and arguments of pericenter $\omega_{i}$ are osculating parameters valid for the reference epoch $t_{0}=2454422.79$ BJD.
Figure 1: Summary of RV fitting results and uncertainties from the MCMC posterior distributions for HD 45364, using data from the pre- and post-upgrade HARPS (HARPS1 and HARPS2, respectively) and the HIRES instruments. The top panel shows our fit RV signal plotted over the observed RVs, with $1\sigma$, $2\sigma$ and $3\sigma$ uncertainties illustrated by the red shaded regions. The maximum log-likelihood solution is plotted as the green line. The HARPS data points to the right of the black line are new data obtained from the HARPS-RVBank archive (Trifonov et al., 2020) since the original analysis of Correia et al. (2009). Time is given in units of Barycentric Julian Date (BJD) - $2.4\times 10^{6}$ days. The normalized residuals of the best-fit solution are plotted in the bottom panel.
@@
Figure 2: The marginal $2$D MCMC posterior distributions of $\sin\left(I\right)$ against planet masses $m_{b}$ (left) and $m_{c}$ (right), with the planets assumed to be in coplanar orbits. The contours show the $1\sigma$, $2\sigma$ and $3\sigma$ bounds of the posterior distribution. The points corresponding to the best-fit solution obtained by maximizing the log-likelihood $\ln\mathcal{L}$ as given in Table 1 are marked in red.
@@ 2.2 The influence of model priors and measurement noise on resonant dynamics
The solid lines in Figure 3 show the cumulative distribution for the maximum deviation of the planets’ resonant angles from equilibrium, measured from short $N$-body integrations spanning $500$ orbits of the inner planet for a random selection of $50,000$ samples from our full MCMC posterior. For each posterior sample, the resonant angles $\theta_{b}=3\lambda_{c}-2\lambda_{b}-\varpi_{b}$ and $\theta_{c}=3\lambda_{c}-2\lambda_{b}-\varpi_{c}$ were recorded at $5000$ uniformly spaced times in the integration and the maximum deviations from their respective equilibrium values of 00 and $\pi$ were computed. Figure 3 clearly indicates that most of the posterior samples have large libration amplitudes, with roughly half of the posterior samples showing libration amplitudes greater than $90^{\circ}$ in both resonant angles and a majority of posterior samples ($\gtrsim 60\%$) showing circulation in angle $\theta_{c}$. These results nominally support the conclusion of Li et al. (2022) that the HD 45364 planets are not fully locked in the 3:2 MMR. [^4] Strictly speaking, the circulation of $\theta_{c}$ does not preclude the system’s residence in the 3:2 MMR (see, e.g. Petit et al., 2020). However, resonance capture via smooth migration is expected to produce resonant configurations with small libration amplitudes, nominally at odds with the posterior distribution of libration amplitudes.
Figure 3: The cumulative distribution of the maximum deviation from equilibrium for the resonant angles $\theta_{b}$ (blue) and $\theta_{c}$ (orange), measured over short $N$-body integrations, of $50,000$ samples drawn randomly from the MCMC posteriors for the $N$-body dynamical model fit under uninformative priors described in Section 2.1 (solid lines) and the model described in Section 2.2 fit under a resonance-weighted prior (i.e. weighted toward small libration amplitudes), with $S=0.1$ (dashed lines). A large fraction of orbital configurations in the uninformative prior model’s MCMC posteriors circulate or have large libration amplitudes ($>90\degree$). In contrast, virtually all resonant angles in the resonance-weighted prior model’s MCMC posteriors have small libration amplitudes ($<90\degree$), showing that the chosen model priors strongly affect the inferred libration amplitudes of the system.
However, Jensen & Millholland (2022) demonstrate that measurement noise can cause a systematic bias to resonant libration amplitudes inferred from RVs. This bias is essentially caused by the fact that higher libration amplitude dynamical configurations occupy a much larger phase-space volume. Consequently, under model priors that are relatively uniform over phase space, noisier data will lead to broader posterior distributions that, in turn, become increasingly skewed towards large libration amplitudes or circulating resonant angles. In this section, we explore the sensitivity of the inferred resonant libration amplitudes of the HD 45364 planets to assumptions about measurement noise and model priors.
@@
where $\mathcal{L}$ is the likelihood function of our uninformative prior RV model given in Equation (2) and $\hat{\sigma}_{\mathrm{jit},j}$ are the best-fit jitter values determined via a maximum-likelihood fit to the RV data using the uninformative prior RV model. In Figure 4, we numerically maximize the modified log-likelihood, $\ln\mathcal{L}^{\prime}$, on a grid of fixed values for the parameters $f$ and $S$. We then plot contour levels of $\ln\mathcal{L}$, i.e., the unmodified, uninformative RV model log-likelihood, evaluated at the numerically determined parameters that maximize the modified likelihood, $\mathcal{L}^{\prime}$. The $n\sigma$ confidence regions are estimated as $\Delta\ln\mathcal{L}=-\frac{1}{2}n^{2}$, where $\Delta\ln\mathcal{L}$ is the difference between the log-likelihood function and its global maximum value. If the data strongly favored dynamical configurations in which the resonant angles circulated or possessed large libration amplitudes, we would expect to find that the values of $\ln\mathcal{L}$ decrease significantly for sufficiently small values of $S$. However, this is not what we see in Figure 4. Instead, contours of constant likelihood are nearly independent of the assumed value of $S$ except at the lowest values of the jitter scaling factor, $f$. Thus, we conclude that the prevalence of dynamical configurations with circulating and large libration amplitude resonant angles found among the posterior samples of our uninformative prior RV model is the result of these dynamical configurations occupying larger prior volume, rather than the data disfavoring low libration amplitude configurations. Furthermore, the fact that low libration amplitude solutions become disfavored relative to higher amplitude solutions for small values of $f$ indicates that underestimated measurement noise can lead one to improperly infer large libration amplitudes.